OpenStudy (anonymous):

Let a and b be square numbers. Is gcd(a,b) also square? Prove your answer, or find a counterexample. (Hint: First show that a number is square if and only if its prime decomposition contains only even exponents.) _i solved the hint~ but have no idea of next step~

OpenStudy (anonymous):

1,4,9,16,25 etc.... these are square numbers because 1^2 = 1, 2^2 = 4, 3^2 = 9 4^2 = 16 and 5^2 = 25 all with even exponents

OpenStudy (anonymous):

100 and 4. gcd is 2, 4 and 9. gcd is 1. A lot of other examples to prove..gcd of two square numbers is not always a square number

OpenStudy (anonymous):

dont google that~ the statement is actually true(can be proved), btw gcd(100,4)=4.

ganeshie8 (ganeshie8):

another hint : square root of gcd exists in both the numbers :)

OpenStudy (anonymous):

yea that makes sense, but i dont know how. omg, so tricky '

ganeshie8 (ganeshie8):

Okay. heres the proof : since a & b are square numbers, they can be written as even powers of prime nuumbers : p,q,r,t let : \(\huge a = p^{2}q^{8}r^{6} \\ \huge b = p^2t^4r^4\)

ganeshie8 (ganeshie8):

@satellite73 :)

OpenStudy (anonymous):

gcd is a square because gcd itself can be factored as the product of primes say \(p_1^{\alpha_1}\times p_2^{\alpha_2}\times ...\times p_n^{\alpha_n}\) but if \(p_i|a\) then so does \(p_i^{2k}\) for some \(k\) since you have proved that all the exponents are even similarly \(p_i^{2j}\) divides \(b\) and since \(p_i^{\alpha_i}\) divides both \(a\) and \(b\) for each \(i\) you have \(\alpha_i=2l\) for some \(l\)

OpenStudy (anonymous):

thx u people for the hints~ helps a lot. i think i got the idea :D

OpenStudy (beginnersmind):

I wonder if we can prove this without referring to prime factorization. It's true that \[\gcd(a^{2},b^{2}) = (\gcd(a,b))^2\] but how do you prove it? Let's write a = c * gcd(a,b) b = d * gcd(a,b) with gcd(c,d) = 1 so we have \[a^{2} = c^{2} * (gcd(a,b)^{2}\] \[b^{2} = d^{2} * (gcd(a,b)^{2}\] \[gcd(c^{2},d^{2}) = 1\] So on the one hand, we have \[(gcd(a,b))^{2} | a \] \[(gcd(a,b))^{2} | b \] proving that it's a common divisor. On the other hand if k is a divisor of a, b and \(gcd(a,b)^2\) it is of the form \(k*gcd(a,b)^2\) where k divides both \(c^2\) and \(d^2\). So k =1 proving that \[\gcd(a^{2},b^{2}) = (\gcd(a,b))^2\]

OpenStudy (anonymous):

wow~ thank you for this~ :) @beginnersmind